OpenStudy (toxicsugar22):

At a concert By the Who in 1976, the sound level 50 meters from the stag registered 120 decibles. The treshold of pain for the human ear is 90 decibles. Calculate the sound intensity of the Who concert and the sound intenisity of the treshold of pain

OpenStudy (heisenberg):

I think you need to combine the sound intensity level formula and the inverse square law for this guy.

OpenStudy (toxicsugar22):

WILL YOU HELP ME DO THAT

OpenStudy (anonymous):

why inverse square law?

OpenStudy (toxicsugar22):

CAN YOU HELP ME

OpenStudy (toxicsugar22):

PLEASE

OpenStudy (toxicsugar22):

YEAH THANKS BUT i DONT KNOW HOW TO START IT

OpenStudy (toxicsugar22):

WILL YOU HELP ME STAT=RT THIS PROBLEM PLEASE IM SO STRESSED

OpenStudy (anonymous):

please read it carefully

OpenStudy (toxicsugar22):

I DID

OpenStudy (toxicsugar22):

PLEASE i WILL BE GRATEFULL TO YOU

OpenStudy (toxicsugar22):

WILL YOU GO STEP BY STEP WITH ME TO OLVE THIS

OpenStudy (heisenberg):

I will try to help, but I am not too sure. But please, take caps lock off.

OpenStudy (toxicsugar22):

ok

OpenStudy (heisenberg):

\[Intensity = \frac{Power}{4*\pi*r^2}\]

OpenStudy (heisenberg):

sorry, not 'power' but 'sound wave pressure'

OpenStudy (toxicsugar22):

ok

OpenStudy (heisenberg):

to find P, use this: \[P = P_r * 10^\frac{SPL}{20}\] P_r is the reference point

OpenStudy (heisenberg):

SPL = decibel level

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

will you help further

OpenStudy (toxicsugar22):

so i understand

OpenStudy (toxicsugar22):

fr next time

OpenStudy (heisenberg):

What don't you understand?

OpenStudy (toxicsugar22):

calculaed the intensity of the concert and the pain

OpenStudy (toxicsugar22):

can you help

OpenStudy (heisenberg):

yes you can use those two formulas. First calculate the pressure, then you can use that to calculate the intensity.

OpenStudy (toxicsugar22):

yes can u supervise me

OpenStudy (toxicsugar22):

help

OpenStudy (toxicsugar22):

so I plug in 120 for SLP

OpenStudy (heisenberg):

Yes

OpenStudy (heisenberg):

according to wikipedia, P_ref is often regarded as 20 µPa in air, so I would use that for P_r

OpenStudy (toxicsugar22):

SO IT IS P=20*10(120/20)

OpenStudy (toxicsugar22):

IS IT RIGHT

OpenStudy (heisenberg):

yes that'll give you the pressure. then you need to find the intensity from a distance.

OpenStudy (heisenberg):

glad the caps are back.

OpenStudy (toxicsugar22):

so i did 20*10 which is 1200

OpenStudy (toxicsugar22):

then i did 120/20 which is 6

OpenStudy (heisenberg):

no, there is an exponent there: \[20*10^\frac{120}{20}\]

OpenStudy (heisenberg):

= 20 000 000

OpenStudy (toxicsugar22):

i did that and i DID iT WRONG

OpenStudy (toxicsugar22):

like i have this website from school and it checks my work and I did it wrong

OpenStudy (toxicsugar22):

can u guide me

OpenStudy (toxicsugar22):

again

OpenStudy (toxicsugar22):

srry to bother you

OpenStudy (toxicsugar22):

heisenberg you there

OpenStudy (heisenberg):

what did you get? I get = 636.619772

OpenStudy (toxicsugar22):

i got that also

OpenStudy (toxicsugar22):

it not right

OpenStudy (toxicsugar22):

tAT IS NOT RIGHT HEISENBERG

OpenStudy (toxicsugar22):

please help again

OpenStudy (toxicsugar22):

nm